### v1

# Nearest Neighbor CCP-Based Molecular Sequence Analysis

Abstract
Molecular sequence analysis is crucial for comprehending several biological processes, including protein-protein interactions, functional annotation, and disease classification. The large number of sequences and the inherently complicated nature of protein structures make it challenging to analyze such data. Finding patterns and enhancing subsequent research requires the use of dimensionality reduction and feature selection approaches. Recently, a method called Correlated Clustering and Projection (CCP) has been proposed as an effective method for biological sequencing data. The CCP technique is still costly to compute even though it is effective for sequence visualization. Furthermore, its utility for classifying molecular sequences is still uncertain. To solve these two problems, we present a Nearest Neighbor Correlated Clustering and Projection (CCP-NN)-based technique for efficiently preprocessing molecular sequence data. To group related molecular sequences and produce representative supersequences, CCP makes use of sequence-to-sequence correlations. As opposed to conventional methods, CCP doesn’t rely on matrix diagonalization, therefore it can be applied to a range of machine-learning problems. We estimate the density map and compute the correlation using a nearest-neighbor search technique. We performed molecular sequence classification using CCP and CCP-NN representations to assess the efficacy of our proposed approach. Our findings show that CCP-NN considerably improves classification task accuracy as well as significantly outperforms CCP in terms of computational runtime.
Index Terms:
Nucleotides, CCP, Spike sequence, Dimensionality Reduction
I Introduction
Molecular sequences are a crucial part of the dynamic changes in sequence composition that control biological processes. Researchers have analyzed these molecular sequences and made progress toward a better understanding of physiology, biological development, and disease [1]. To understand various biological mechanisms and disorders, it is essential to understand how proteins interact with one another and carry out certain functions [2]. Unfortunately, because of their vastness, complexity, and lack of distinct patterns, it is still difficult to analyze thousands of molecular sequences at once. The development of medicines and treatments for human diseases is hampered by these obstacles. Deciphering the functions that proteins serve in various physiological and pathological circumstances is the goal of the field of proteomics [3]. Large sets of molecular sequences may now be made feasible by recent technical advancements to be analyzed to find patterns that could ultimately result in the creation of brand-new medications and vaccines [4]. However, processing such massive amounts of data necessitates the use of cutting-edge computing tools and statistical techniques to extract pertinent information.
Additionally, It is well known that the data with a highly dimensional feature space will become sparse, which makes it difficult for statistical analysis to identify statistical significance and key factors. To facilitate prediction, analysis, and visualization, it is, therefore, preferable to minimize the dimensionality of high-dimensional data. Due to these challenges, a variety of dimensionality reduction (DR) techniques has been developed that can accurately reflect the inherent correlations in the original data on a low-dimensional space. There are several linear and non-linear DR methods have been proposed such as principal component analysis (PCA), Linear discriminant analysis (LDA), Multidimensional Scaling (MDS) [5], LargeVis [6] are a few linear dimensionality reduction methods. Whereas kernel PCA [7], Sammon mapping [8] and spectral embedding [9] are non-linear dimensionality reduction techniques.
Correlated Clustering and Projection (CCP), a non-linear dimensionality reduction approach, computes the pairwise correlation matrix of samples and imposes a cutoff distance to prevent the global summation during the projection to increase computational efficiency [10, 11]. CCP has several benefits to offer, such as handling the dimensionality reduction of high sample sizes (because it avoids matrix diagonalization and instead solves a matrix to lower the dimensionality), employing statistical metrics like covariances to quantify the high-level dependence between random feature vectors [12], and it can be used in conjunction with a frequency-domain method for secondary dimensionality reduction to improve the preservation of data’s global structures and increase accuracy [13].
This research is focused on creating a pipeline for analyzing molecular sequences using a method based on Nearest neighbor CCP-NN to preprocess the data and produce condensed representations that accurately reflect the original sequences. Here instead of computing the distance between data points, we use Nearest Neighbor (NN) to compute the nearest neighbor distance. We propose a method we name CCP-NN which is based on CCP to give an extra edge over the original CCP. The main contributions of our research are as follows:
  1. 1.
    We propose a novel approach based on CCP to preprocess molecular sequence data, which leverages sequence-sequence correlations to generate representative super-sequences.
  2. 2.
    We demonstrate the effectiveness of CCP and CCP-NN by evaluating their performance in molecular sequence classification tasks. Our results indicate that CCP-NN significantly enhances the accuracy of classification compared to other methods.
  3. 3.
    We offer a thorough framework for examining molecular sequence data by utilizing the capabilities of CCP. Our method facilitates effective feature selection, dimensionality reduction, and visualization.
II Related Work
Sequence classification is a well-researched issue in bioinformatics [14]. A phylogenetic approach is frequently used in more conventional techniques of analyzing sequencing data [15], however, they are not scalable due to higher computational cost. To counter the issue, some machine learning (ML) approaches, including alignment-based [16, 17] and alignment-free [18] embedding approaches have become popular for ML tasks such as classification and clustering. Due to the extremely high dimensionality of the feature vector, these techniques do, however, also have scalability issues. The classification of biological sequences also makes use of the kernel matrix technique [19]. The Wasserstein distance (WD) is used in [20] to extract the features. Some efforts have been made to improve computational performance, such as Locality Sensitive Hashing (LSH) [21], which can train models faster and more accurately. The hash function in [22] is used to generate an approximate word embedding for language processing. However, collisions might occur in the resulting vectors, which reduces the embedding’s effectiveness. The use of bloom filters for mistake correction in raw read data to aid a de novo assembly was demonstrated by authors in [23]. However, these methods are prone to cause loss of information. Some of the most popular methods are principal component analysis (PCA) [24], Multidimensional Scaling [5], and LargeVis [6]. The curse of dimensionality and difficulties with the analysis of outliers are another issue [25]. When it comes to data noise, missing data, and poor-quality data, it can be quite unstable [25]. Sequence analysis is still difficult to perform despite extensive effort because of the high dimensionality and quantity of data [26]. Therefore, methods that can manage the dimensionality reduction of high sample numbers and do not involve matrix diagonalization are required.
III Proposed Approach
We divide this section into two parts where we first discuss the original Correlated Clustering and Projection (CCP) method proposed by [10, 11]. After that, we discuss the proposed method, called Nearest Neighbor CCP-NN in detail.
III-A Correlated Clustering and Projection (CCP) Algorithm
The Correlated Clustering and Projection (CCP) algorithm [10, 11] is a data clustering and dimensionality reduction technique that identifies and groups correlated features within a high-dimensional dataset. The algorithm operates by partitioning the features into clusters based on their correlation patterns and then projecting the data onto the subspace spanned by the identified clusters. The purpose of this algorithm is to capture the underlying structure of the data by focusing on feature subsets that exhibit strong correlations, thereby facilitating meaningful analysis and visualization. Given a dataset with $N$ samples and $M$ features represented by the matrix $\mathbf{X}$, the CCP proceeds as follows:
Step 1 (Data Preprocessing)
The algorithm begins by calculating the variance of each feature to identify non-zero variance features, which are essential for meaningful clustering.
Step 2 (Selecting Features for Clustering)
The next step is to select a subset of features for clustering based on the variance. The algorithm chooses the top $numCutoff$ features, which is a user-defined parameter representing the percentage of non-zero variance features to retain.
Step 3 (K-Means Clustering)
The selected features are clustered using the K-Means algorithm with $n\_components-1$ clusters, where $n\_components$ (a hyperparameter) is the desired number of clusters.
Step 4 (Partitioning Non-Clustered Features)
The features that were not assigned to any cluster due to low variance are grouped into a separate cluster. This makes a new cluster that contains the remaining features.
Step 5 (Computing Density Map)
For each cluster, the algorithm computes a density map to capture the correlation between features within the cluster (see Algorithm 1). The density map is estimated using either an exponential kernel or a Lorentz kernel, which are defined as follows:
$\text{Exponential Kernel}=>K(x)=e^{-\left(\frac{x}{\text{scale}}\right)^{\text{power}}}$ (1)
$\text{Lorentz Kernel}=>K(x)=\frac{1}{1+\left(\frac{x}{\text{scale}}\right)^{\text{power}}}$ (2)
where $x$ represents the pairwise distance between two features, scale is a scaling factor, and power is a user-defined parameter.
In Algorithm 1, the CCP performs several steps to compute the correlation and estimate the density map based on the given inputs and parameters. It begins by calculating the pairwise distances between the selected features. If the transformation flag is set to true, it calculates the distances between the features in the input data and the reference data. Otherwise, it calculates the distances between the features in the reference data. Next, if the scaling factor is not already calculated for the specified component, it proceeds to compute the average minimum distance, which is important for scaling the density estimation. Similarly, if the cutoff value is not already set for the specified component, it computes the average and standard deviation (SD) of the pairwise distances. The cutoff is then defined as the average plus three times the SD. It helps determine the threshold beyond which correlation values are considered negligible. The algorithm calculates the scaling factor by multiplying the user-defined scaling parameter with the previously computed average minimum distance. The scaling factor is used to adjust the scale of the density estimation. Finally, the algorithm estimates the density map, also known as the correlation, based on the calculated scaling factor and cutoff value. This estimation is done by computing the density of the pairwise distances using a density estimation function. The resulting density map represents the correlation between the selected features.
1:  computeCorrelations($index\_component$, $index\_Feat$, $X$, $transform$)
2:  if $transform$ then
3:     corr $\leftarrow$ pairwise_distances(X[:, index_Feat], self.X[:, index_Feat], self.metric)
4:  else
[…]
6:  end if
[…]
9:  end if
[…]
11:     $avg\leftarrow\textsc{mean}(corr)$
12:     $std\leftarrow\textsc{std}(corr)$
[…]
14:  end if
15:  $Scale\leftarrow\text{self.scale}\times\text{self.avgmindist}[idx\_comp]$
16:  cutOffVal $\leftarrow$ self.cutoff[ind_comp]
17:  $corr\leftarrow\textsc{computeDensity}(corr,Scale,cuttOffVal)$
18:  Return $corr$
Algorithm 1 Pseudocode for computing Correlation for CCP.
Step 6 (Density-based Clustering)
The density map obtained for each cluster is used for a final density-based clustering. Features are assigned to clusters based on their density values, where higher density indicates a stronger correlation.
Step 7 (Projection)
Finally, the data is projected onto the subspace spanned by the identified clusters. Each sample is represented as a vector of density values corresponding to each cluster. This new projected representation into the subspace, called $\phi_{CCP}$ is used as the low dimensional embedding representation for the given data point.
Remark 1
For a detailed description of the original CCP algorithm, readers are referred to [10, 11].
III-B Nearest Neighbors Based CCP
In the nearest neighbor (NN) version of CCP (our proposed method), all steps from 1 to 7 are followed from the original CCP as described in the above subsection. The main modification is made in Step 5, where we compute the density map using the NN algorithm for efficient and fast computation of the density map. The pseudocode for computing the density map is given in Algorithm 2, where the NearestNeighborComputeCorr function incorporates the use of an NN search technique, specifically the AnnoyIndex data structure [27], to calculate the correlation and estimate the density map. The steps involved in this process are as follows: An AnnoyIndex is created with the specified metric, and the features from the reference data are added to the index. The function checks if the transformation flag is set to true then we find a correlation by vector else if it is set to false, we find a correlation by item.
1:  NearestNeighborComputeCorr$(idx\_component,idx\_Feat,X,transform)$
[…]
4:     $idx.add\_item(i,\text{self.X}[:,idx\_Feat][I])$
5:  end for
6:  $idx.build(-1)$
7:  if $transform$ then
8:     $corr\leftarrow[\textsc{idx.get\_nns\_by\_vector}(\text{Feat},1)\text{for Feat in }X[:,idx\_Feat]]$
9:  else
10:     $corr\leftarrow[\textsc{idx.get\_nns\_by\_item}(i,1)\text{ for }i\text{ in range}(\text{len}(\text{self.X}[:,idx\_Feat]))]$
11:  end if
12:  $corr\leftarrow\textsc{reshape}(corr,(-1,1))$
[…]
15:  end if
[…]
17:     $avg\leftarrow\textsc{mean}(corr)$
18:     $std\leftarrow\textsc{std}(corr)$
[…]
20:  end if
21:  $Scale\leftarrow\text{self.scale}\times\text{self.avgmindist}[idx\_component]$
22:  cuttOffVal $\leftarrow$ self.cutoff[idx_component]
23:  $corr\leftarrow\textsc{computeDensity}(corr,Scale,cuttOffVal)$
24:  return $corr$
Algorithm 2 Pseudocode for computing correlation for Nearest Neighbor CCP.
Remark 2
In NN-CCP, we utilize the AnnoyIndex data structure, which is an efficient implementation of NN. Instead of computing the exact distances between data points, AnnoyIndex builds an index that allows for fast retrieval of NN. This significantly speeds up the computation of pairwise distances, making it useful for large datasets and high-dimensional spaces.
Once the AnnoyIndex is constructed, the function retrieves the NN for each feature in the input data. If the transformation flag is true, it retrieves the NN based on the features in the input data. If the flag is false, it retrieves the NN based on the features in the reference data. The retrieved nearest neighbors (NN) are reshaped into a proper format for further processing. Similar to Algorithm 1, the function checks if the scaling factor needs to be computed for the specified component. If the scaling factor is not already calculated, it proceeds to compute the average minimum distance, which is crucial for scaling the density estimation.
If the cutoff value is not set, it computes the average and SD of the correlation values obtained from the NN. The cutoff is then defined as the average plus three times the SD. This cutoff value helps determine the threshold beyond which correlation values are considered negligible. The scaling factor is calculated by multiplying the user-defined scaling parameter with the previously computed average minimum distance. This scaling factor is used to adjust the scale of the density estimation.
Finally, the function estimates the density map, also known as the correlation, by applying the density estimation function to the correlation values obtained from the NN. This density map represents the correlation between the selected features, taking into account the scaling factor and the cutoff value. In summary, Algorithm 2 incorporates the use of NN to calculate the correlation and estimate the density map. It involves constructing an AnnoyIndex, retrieving the NN, reshaping the obtained correlations, computing the scaling factor and cutoff value, and estimating the density map based on these values. This approach allows for efficient computation of correlations and density estimation, particularly for high-dimensional data.
After computing the density map, Step 6 and Step 7 are followed similarly to the original CCP as described above (i.e., used to compute $\phi_{CCP}$). After Step 7, we get the new projected representation into the subspace, called $\phi_{CCP\_NN}$ (where NN stands for Nearest Neighbor), which is used as the low dimensional embedding representation for the given data point.
III-C Algorithm Complexity
III-C1 CCP
To begin, the computation of variance along each feature axis takes $\mathcal{O}(N\cdot M)$, where $N$ is the number of samples and $M$ is the number of features. Sorting the variance values to select the top $f\ (\text{where}\ f\leq M)$ features takes $\mathcal{O}(f\log f)$. Following this, the K-Means clustering step typically depends on the number of iterations $n_{\text{iter}}$, the number of features $f$, the number of clusters $n_{c}$, and the number of samples $N$. This step has a time complexity of $\mathcal{O}(n_{c}\cdot f\cdot n_{\text{iter}}\cdot N)$.
To this end, the overall time complexity of the algorithm can be expressed as:
$\mathcal{O}(N\cdot M+f\log f+n_{c}\cdot f\cdot n_{\text{iter}}\cdot N)$ (3)
Since the K-Means setup dominates the variance computation and sorting steps, this simplifies to:
$\mathcal{O}(n_{c}\cdot n_{\text{iter}}\cdot f\cdot N)$ (4)
The total space complexity is dominated by the size of the input matrix and the memory used for K-Means clustering. Therefore, the overall space complexity is:
$\mathcal{O}(N\cdot M)$ (5)
Furthermore, in order to compute the density map to capture the correlation between features within a cluster, pairwise distance calculation is required, which is the most computationally expensive operation. If $f_{i}$ is the number of features selected for the $i$-th cluster, the pairwise distance calculation takes $\mathcal{O}(N^{2}\cdot f_{i})$. This is because pairwise distance calculation compares each of the $N$ samples with every other sample across $f_{i}$ features. Summing over all $n_{c}$ components, the total time complexity becomes:
$\mathcal{O}\left(\sum_{i=1}^{n_{c}}N^{2}\cdot f_{i}\right)=\mathcal{O}(N^{2}\cdot f)$ (6)
where $f=\sum_{i=1}^{n_{c}}f_{i}$ is the total number of features across all components. The space complexity associated with this procedure is $\mathcal{O}(N^{2})$. Therefore, the overall complexity is;
$\mathcal{O}\big{(}N\cdot f(n_{c}\cdot n_{iter}+N)\big{)}$ (7)
and the space complexity is;
$\mathcal{O}\big{(}N(M+N)\big{)}$ (8)
III-C2 CCP-NN
On the other hand, if we were to use Approximate Nearest Neighbor (ANN) via Annoy Index as a proxy for pairwise distance computation, we would significantly benefit in terms of computation time. Building an Annoy Index takes $\mathcal{O}(N\log N\cdot f)$, because the algorithm builds a forest of random projection trees. Querying the Annoy Index for nearest neighbors is approximately $\mathcal{O}(\log N)$ per query, and since this is repeated for all $N$ samples, the complexity for querying is $\mathcal{O}(N\log N)$.
With $n_{c}$ total clusters, the total time complexity becomes
$\mathcal{O}(n_{c}\cdot n_{\text{iter}}\cdot f\cdot N)+\mathcal{O}\big{(}\sum_{i=1}^{n_{c}}N\log N\cdot f_{i}\big{)}=\mathcal{O}\big{(}N\cdot f(n_{c}\cdot n_{iter}+\log N)\big{)}$ (9)
and the space complexity is;
$\mathcal{O}\big{(}N(M+\log N\cdot f)\big{)}$ (10)
Clearly, CCP-NN has an advantage over CCP in terms of speed and memory requirement.
III-D Convergence Analysis
The key step in CCP-NN is the estimation of the density map, which is used to capture the correlation between features within a cluster. This is done using the nearest neighbor search, especially using the Annoy Index data structure [27].
Let $X\in\mathbb{R}^{N\times M}$ be a dataset with $N$ samples and $M$ features (dimension). Given a set of features $\{x_{i}\}_{i=1}^{M}$ within a cluster, the density estimate at each point $x_{i}$ is determined by the proximity of its nearest neighbors. The Annoy Index facilitates the retrieval of the nearest neighbors, denoted by $\mathcal{N}_{k}(x_{i})$, where $k$ is the number of neighbors considered.
Assume the following;
  • $X\sim\mathbb{P}(X)$ with a well defined density function $p(x)$
  • Nearest neighbor search in CCP-NN provides a close approximation to the true nearest neighbors, with an error margin $\epsilon$
Now, let $\tilde{\mathcal{N}}_{k}(x_{i})$ represent the nearest neighbors returned by the Annoy Index, and let $\mathcal{N}_{k}$ represent the true nearest neighbors. The accuracy of the Annoy Index guarantees that:
$\mathbb{E}[||\mathcal{N}_{k}(x_{i})-\tilde{\mathcal{N}}_{k}(x_{i})||]\leq\epsilon$ (11)
where $\epsilon>0$ depends on the dimensionality of the data and the parameters of the Annoy Index.
Given the convergence of the nearest neighbor search, we can now analyze the consistency of the density estimation process. Let $\hat{p}(x_{i})$ be the density estimate at $x_{i}$ obtained by CCP-NN, and let $p(x_{i})$ be the true density. The density estimate is given by;
$\hat{p}(x_{i})=\frac{1}{hk}\sum_{x_{j}\in\hat{\mathcal{N}}_{k}(x_{i})}K\bigg{(}\frac{x_{i}-x_{j}}{h}\bigg{)}$ (12)
where $K$ is a kernel function and $h$ is a bandwidth parameter. Using Triangle-Inequality:
$|\hat{p}(x_{i})-p(x_{i})|\leq|\hat{p}(x_{i})-p(x_{i},\hat{N}_{k}(x_{i}))|+|p(x_{i},\hat{N}_{k}(x_{i}))-p(x_{i})|$ (13)
where $p(x_{i},\hat{N}_{k}(x_{i}))$ denotes the density estimate using the true nearest-neighbor. We take expectation on both sides and the first term of R.H.S in the inequality above can be bounded by $\mathcal{O}(\epsilon)$ due to the accuracy of the Annoy Index. The second term can be bounded using standard kernel density estimation convergence results [28].
Therefore, the upper bound on the error estimate is:
$\mathbb{E}\big{[}|\hat{p}(x_{i})-p(x_{i})|\big{]}\leq\mathcal{O}(\epsilon+h^{4}+1/kh)$ (14)
IV Experimental Setup
In this section, we describe the different datasets used for experiments. We also go through the baseline methods and evaluation metrics we use for the classification. A Windows 10 $64$-bit machine with an Intel(R) Core i5 processor operating at $2.10$ GHz and $32$ GB of memory is used for all experiments. Our pre-processed datasets and code are available online for reproducibility 111available in the published version.
We use three datasets in this study to assess the effectiveness of the proposed method. We employ t-distributed stochastic neighbor embedding (t-SNE) [29] to examine the data for any natural (hidden) grouping. The t-SNE plots for various embedding techniques are displayed in Figure 1,  2, and  3 for Protein Subcellular, Coronavirus Host, and Human DNA datasets, respectively. We can observe that t-SNE can group similar classes in the case of Autoencoder with CCP-NN.
(a) OHE (CCP)
(b) Spike2Vec (CCP)
(c) PWM2Vec (CCP)
(d) Autoencoder (CCP)
(e) OHE (CCP-NN)
(f) Spike2Vec CCP-NN
(g) PWM CCP-NN
(h) Auto-En. CCP-NN
Figure 1: t-SNE plots (Protein Subcellular Data) for different structure embeddings and Clustering and Projection methods (CCP and CCP-NN). The figure is best seen in color.
(a) OHE (CCP)
(b) Spike2Vec (CCP)
(c) PWM2Vec (CCP)
(d) Autoencoder (CCP)
(e) OHE (CCP-NN)
(f) Spike2Vec CCP-NN
(g) PWM CCP-NN
(h) Auto-En. CCP-NN
Figure 2: t-SNE plots (Coronavirus Host Data) for different structure embeddings and Clustering and Projection methods (CCP and CCP-NN). The figure is best seen in color.
(a) OHE (CCP)
(b) Spike2Vec (CCP)
(c) PWM2Vec (CCP)
(d) Autoencoder (CCP)
(e) OHE (CCP-NN)
(f) Spike2Vec CCP-NN
(g) PWM CCP-NN
(h) Auto-En. CCP-NN
Figure 3: t-SNE plots (Coronavirus Host Data) for different structure embeddings and Clustering and Projection methods (CCP and CCP-NN). The figure is best seen in color.
To classify the molecular sequences, we employed several ML models, including Support Vector Machine (SVM), Naive Bayes (NB), Multi-Layer Perceptron (MLP), K-Nearest Neighbors (KNN), Random Forest (RF), Logistic Regression (LR), and Decision Tree (DT). We evaluated classification performance using average accuracy, precision, recall, F1 (weighted), F1 (macro), Receiver Operator Characteristic Curve (ROC) Area Under the Curve (AUC), and training runtime. To preserve the original data distribution, the data for each classification task is divided into 60-10-30% train-validation-test sets using stratified sampling. To obtain more consistent findings, we also conduct our tests by averaging the performance outcomes of $5$ runs. We carefully considered baselines from several embedding generation categories, including feature engineering, conventional kernel matrix generation, neural networks, pre-trained language models, and pre-trained transformers for protein sequences. Table I contains the specifics for the baseline models.
Method Category Detail Source
OHE Feature Engineering The numerical vector is generated with simple one-hot encoding for each amino acid in the sequence.  [16]
Spike2Vec Uses the sliding window (of size k) to get $k$-mers and its count in the sequence to generate feature vectors.  [18]
PWM2Vec Uses the concept of the position-weight matrix (PWM) to generate embeddings  [17]
String Kernel Kernel Matrix Designs $n\times n$ kernel matrix that can be used with kernel classifiers or with kernel PCA to get feature vector.  [19]
WDGRL Neural Network (NN) Take one-hot representation of biological sequence as input and design NN-based embedding method by minimizing loss  [20]
AutoEncoder  [30]
SeqVec Pretrained Large Language Model (LLM) Takes biological sequences as input and fine-tunes the weights based on a pre-trained model to get final embedding.  [31]
ProteinBERT Pretrained Transformer A pre-trained protein sequence model to classify the given biological sequence using Transformer/Bert  [32]
TAPE Pretrained Transformer A LLM model with a self-supervised pretraining method for molecular sequence embedding generation.  [33]
TABLE I: Baseline methods.
IV-A Baseline Methods
Among the baseline methods discussed in Table I, we selected $4$ popular embedding generation models, including One Hot Encoding (OHE) [16], Spike2Vec [18], PWM2Vec [17], and Autoencoder [30] to be used as input to both vanilla CCP and CCP-NN for dimensionality reduction. These methods were simple to use (in terms of implementation compared to complex models like SeqVec, TAPE, and Protein Bert), easy and fast to compute (compared to WDGRL, which is computationally expensive and takes long computational time), and generate embeddings directly, which can be used for dimensionality reduction (unlike String kernel, which generates a kernel matrix, which has to be converted to embeddings using kernel PCA, which could cause loss of information).
IV-B Dataset Statistics
We use $3$ datasets for our experiments. The datasets employed are listed below:
Protein Subcellular Locations The dataset Protein Subcellular Locations Dataset we employ consists of $5959$ unaligned protein sequences, each corresponding to a different subcellular location [34]. The labels for the classification task are these subcellular sites and there are $11$ unique labels in our dataset. These labels correspond to the proteins of plant cells and fungal cells, while animal cells share all localizations with them. Table II provides the distribution of classes.
Subcellular Locations No. of Sequences Subcellular Locations No. of Sequences
Cytoplasm 1411 Endoplasmic Reticulum 198
Plasma Membrane 1238 Peroxisome 157
Extracellular Space 843 Golgi Apparatus 150
Nucleus 837 Lysosomal 103
Mitochondrion 510 Vacuole 63
Chloroplast 449 - -
- - Total 5959
TABLE II: The distribution of sequences in the Protein Subcellular locations data among the subcellular locations.
Coronavirus Host The NIAD Virus Pathogen Database and Analysis Resource(ViPR) [35] and GISAID [36] are used to retrieve the Spike molecular sequences of CoVs for all of the hosts. Table III (in the supplementary material) comprises details about the $21$ host types with their counts of sequences that we gathered through the annotation of the total of $5558$ complete protein sequence.
Host Name # of Sequences Host Name # of Sequences Host Name # of Sequences
Humans 1813 Rats 26 Cats 123
Environment 1034 Pangolins 21 Bovines 88
Weasel 994 Hedgehog 15 Dogs 40
Swine 558 Dolphin 7 Python 2
Birds 374 Equine 5 Monkey 2
Camels 297 Fish 2 Cattle 1
Bats 153 Unknown 2 Turtle 1
- - - - Total 5558
TABLE III: Statistics for Coronavirus Host dataset.
Human DNA The data contains $4380$ unaligned Human DNA nucleotide sequences [37]. A total of $7$ unique labels comprised of a human gene family are G Protein-Coupled, Tyrosine Kinase, Tyrosine Phosphatase, Synthetase, Synthase, Ion Channel, and Transcription Factor. Table IV (in the supplementary material) provides the statistics for the dataset.
Gene Family Num. of Sequences Gene Family Num. of Sequences
G Protein Coupled 531 Tyrosine Kinase 534
Tyrosine Phosphatase 349 Synthetase 672
Synthase 711 Ion Channel 240
Transcription Factor 1343 - -
- - Total 4380
TABLE IV: The distribution of gene family with the count of sequences in the Human DNA data.
V Results And Discussion
In this section, we report classification and runtime results for both CCP and CCP-NN using different datasets and embedding models.
V-A Results For Protein Subcellular dataset
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Protein Subcellular dataset are shown in Table V. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics and achieves a near-perfect predictive classification performance. The performance gain for CCP-NN (using OHE with Decision Tree classifier and using Autoencoder with Decision Tree classifier), compared to the best CCP-based results (Spike2Vec with Random Forest classifier) is $50.3\%$, which highlights a significant improvement in terms of predictive accuracy.
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
CCP ($\phi_{CCP}$) OHE SVM 0.288 0.228 0.288 0.196 0.085 0.517 17.920
NB 0.209 0.445 0.209 0.215 0.171 0.569 0.344
MLP 0.321 0.327 0.321 0.323 0.199 0.563 6.457
KNN 0.238 0.212 0.238 0.197 0.106 0.515 0.288
RF 0.429 0.420 0.429 0.375 0.215 0.580 3.222
LR 0.349 0.274 0.349 0.267 0.127 0.542 7.668
DT 0.306 0.306 0.306 0.306 0.202 0.563 0.653
Spike2Vec SVM 0.393 0.414 0.393 0.380 0.247 0.589 16.626
NB 0.200 0.318 0.200 0.235 0.161 0.549 0.321
MLP 0.360 0.359 0.360 0.358 0.221 0.577 10.421
KNN 0.262 0.281 0.262 0.247 0.147 0.538 0.279
RF 0.495 0.524 0.495 0.440 0.245 0.598 3.292
LR 0.418 0.412 0.418 0.387 0.233 0.582 10.810
DT 0.319 0.318 0.319 0.317 0.203 0.566 1.077
PWM2Vec SVM 0.411 0.424 0.411 0.412 0.316 0.627 11.482
NB 0.224 0.267 0.224 0.221 0.173 0.561 0.284
MLP 0.335 0.335 0.335 0.334 0.214 0.572 7.056
KNN 0.244 0.369 0.244 0.213 0.135 0.524 0.229
RF 0.450 0.511 0.450 0.391 0.214 0.582 2.987
LR 0.445 0.443 0.445 0.429 0.302 0.613 8.984
DT 0.281 0.280 0.281 0.280 0.183 0.553 0.924
Autoencoder SVM 0.300 0.194 0.300 0.226 0.106 0.528 5.469
NB 0.204 0.162 0.204 0.145 0.089 0.535 0.081
MLP 0.236 0.208 0.236 0.217 0.118 0.520 9.564
KNN 0.239 0.224 0.239 0.226 0.133 0.529 0.144
RF 0.298 0.229 0.298 0.235 0.114 0.529 13.596
LR 0.294 0.216 0.294 0.231 0.111 0.529 2.201
DT 0.197 0.198 0.197 0.197 0.114 0.514 2.236
CCP Nearest Neighbor ($\phi_{CCP\_NN}$) OHE SVM 0.463 0.489 0.463 0.472 0.293 0.619 14.200
NB 0.705 0.735 0.705 0.705 0.604 0.791 0.294
MLP 0.569 0.578 0.569 0.571 0.365 0.661 9.380
KNN 0.239 0.362 0.239 0.152 0.071 0.509 0.290
RF 0.939 0.945 0.939 0.931 0.814 0.882 1.976
LR 0.537 0.506 0.537 0.518 0.318 0.634 10.401
DT 0.998 0.998 0.998 0.998 0.995 0.997 0.200
Spike2Vec SVM 0.587 0.597 0.587 0.590 0.410 0.683 8.628
NB 0.400 0.557 0.400 0.416 0.362 0.679 0.231
MLP 0.619 0.625 0.619 0.620 0.422 0.693 6.110
KNN 0.233 0.356 0.233 0.210 0.119 0.524 0.222
RF 0.940 0.945 0.940 0.934 0.840 0.892 1.740
LR 0.615 0.594 0.615 0.591 0.398 0.670 7.800
DT 0.995 0.995 0.995 0.995 0.991 0.995 0.262
PWM2Vec SVM 0.520 0.537 0.520 0.526 0.368 0.659 13.341
NB 0.395 0.521 0.395 0.398 0.342 0.670 0.322
MLP 0.561 0.567 0.561 0.561 0.365 0.661 9.496
KNN 0.198 0.342 0.198 0.133 0.063 0.499 0.261
RF 0.942 0.946 0.942 0.936 0.846 0.895 3.003
LR 0.571 0.551 0.571 0.552 0.369 0.655 11.213
DT 0.995 0.995 0.995 0.995 0.993 0.996 0.452
Autoencoder SVM 0.814 0.671 0.814 0.734 0.411 0.716 0.612
NB 0.950 0.962 0.950 0.952 0.907 0.981 0.065
MLP 0.933 0.911 0.933 0.918 0.861 0.933 2.851
KNN 0.997 0.997 0.997 0.997 0.991 0.996 0.085
RF 0.998 0.998 0.998 0.998 0.994 0.998 2.290
LR 0.439 0.202 0.439 0.274 0.114 0.558 0.815
DT 0.998 0.998 0.998 0.998 0.995 0.998 0.427
TABLE V: Classification results (averaged over $5$ runs) for Protein Subcellular dataset using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
The comparison of the best performing proposed method from Table V (i.e. CCP-NN with Autoencoder) with the existing baseline models (without CCP or CPP-NN) is shown in Table VI. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the proposed method with Autoencoder embedding achieves $28\%$ improvement compared to the second best (i.e. Protein Bert, a pre-trained transformer-based model) and achieves a near-perfect average accuracy score in the case of the Protein Subcellular dataset.
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
OHE SVM 0.530 0.516 0.530 0.509 0.355 0.647 706.196
NB 0.131 0.201 0.131 0.137 0.091 0.514 16.515
MLP 0.390 0.401 0.390 0.390 0.255 0.595 187.259
KNN 0.242 0.114 0.242 0.095 0.036 0.500 8.465
RF 0.404 0.410 0.404 0.329 0.171 0.563 32.347
LR 0.515 0.498 0.515 0.492 0.335 0.637 6.140
DT 0.307 0.300 0.307 0.303 0.193 0.560 25.282
Spike2Vec SVM 0.575 0.579 0.575 0.571 0.483 0.706 111.398
NB 0.253 0.368 0.253 0.253 0.182 0.578 3.095
MLP 0.478 0.489 0.478 0.481 0.345 0.645 36.700
KNN 0.279 0.416 0.279 0.215 0.136 0.531 1.889
RF 0.480 0.528 0.480 0.429 0.237 0.592 7.353
LR 0.564 0.566 0.564 0.555 0.453 0.687 8.075
DT 0.293 0.289 0.293 0.290 0.183 0.554 2.890
PWM2Vec SVM 0.423 0.444 0.423 0.426 0.339 0.640 79.182
NB 0.293 0.312 0.293 0.241 0.206 0.581 0.810
MLP 0.309 0.315 0.309 0.310 0.206 0.568 111.598
KNN 0.285 0.461 0.285 0.247 0.192 0.549 1.964
RF 0.436 0.496 0.436 0.379 0.210 0.577 84.261
LR 0.470 0.476 0.470 0.470 0.351 0.645 96.467
DT 0.306 0.316 0.306 0.310 0.196 0.561 34.803
Autoencoder SVM 0.431 0.447 0.431 0.435 0.315 0.632 95.840
NB 0.228 0.305 0.228 0.205 0.161 0.569 0.316
MLP 0.412 0.389 0.412 0.399 0.253 0.598 126.795
KNN 0.275 0.292 0.275 0.219 0.127 0.529 1.970
RF 0.381 0.347 0.381 0.306 0.163 0.558 30.260
LR 0.464 0.452 0.464 0.455 0.332 0.639 138.959
DT 0.228 0.232 0.228 0.229 0.150 0.533 15.367
String Kernel SVM 0.496 0.510 0.496 0.501 0.395 0.674 5.277
NB 0.301 0.322 0.301 0.265 0.243 0.593 0.136
MLP 0.389 0.390 0.389 0.388 0.246 0.591 7.263
KNN 0.372 0.475 0.372 0.370 0.272 0.586 0.395
RF 0.473 0.497 0.473 0.411 0.218 0.585 7.170
LR 0.528 0.525 0.528 0.525 0.415 0.678 8.194
DT 0.328 0.335 0.328 0.331 0.207 0.568 2.250
WDGRL SVM 0.229 0.098 0.229 0.137 0.057 0.503 1.752
NB 0.206 0.154 0.206 0.158 0.073 0.501 0.008
MLP 0.218 0.136 0.218 0.151 0.067 0.502 11.287
KNN 0.170 0.154 0.170 0.158 0.086 0.500 0.273
RF 0.211 0.167 0.211 0.163 0.079 0.503 2.097
LR 0.229 0.098 0.229 0.137 0.057 0.503 0.112
DT 0.152 0.154 0.152 0.153 0.086 0.498 0.082
SeqVec SVM 0.412 0.425 0.412 0.421 0.306 0.611 10.241
NB 0.205 0.297 0.205 0.196 0.154 0.542 0.125
MLP 0.403 0.377 0.404 0.384 0.231 0.574 21.495
KNN 0.244 0.271 0.245 0.201 0.114 0.511 1.141
RF 0.362 0.323 0.362 0.295 0.155 0.541 5.137
LR 0.451 0.444 0.451 0.421 0.323 0.624 4.427
DT 0.213 0.221 0.213 0.224 0.149 0.517 7.752
Protein Bert _ 0.718 0.715 0.718 0.706 0.572 0.765 16341.85
TAPE SVM 0.637 0.640 0.637 0.636 0.552 0.760 8.553
NB 0.377 0.508 0.377 0.375 0.300 0.662 0.311
MLP 0.590 0.590 0.590 0.589 0.432 0.695 5.296
KNN 0.595 0.600 0.595 0.589 0.468 0.710 0.160
RF 0.600 0.622 0.600 0.572 0.405 0.666 28.819
LR 0.671 0.664 0.671 0.660 0.553 0.746 15.968
DT 0.417 0.424 0.417 0.420 0.300 0.620 11.233
$\phi_{CCP\_NN}$ (ours) - Autoencoder SVM 0.814 0.671 0.814 0.734 0.411 0.716 0.612
NB 0.950 0.962 0.950 0.952 0.907 0.981 0.065
MLP 0.933 0.911 0.933 0.918 0.861 0.933 2.851
KNN 0.997 0.997 0.997 0.997 0.991 0.996 0.085
RF 0.998 0.998 0.998 0.998 0.994 0.998 2.290
LR 0.439 0.202 0.439 0.274 0.114 0.558 0.815
DT 0.998 0.998 0.998 0.998 0.995 0.998 0.427
TABLE VI: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e. CCP-NN with Autoencoder) with baselines on Protein Subcellular dataset. The best value for each embedding is shown underlined. The overall best value for each evaluation metric is shown in bold.
The standard deviation (SD) results (for $5$ runs) for the baselines and the proposed method are shown in Table VII for the Protein Subcellular dataset. We can observe that in the majority of the cases, the SD values are towards the lower end (i.e. $<0.02$), which shows that there is not much variation in the results for different experimental runs having a random train-test split.
Embeddings Algo. Acc. Prec. Recall F1 (Weig.) F1 (Macro) ROC AUC Train Time (sec.)
OHE SVM 0.005339 0.005547 0.005339 0.004908 0.00894 0.003092 3.916014
NB 0.165977 0.047923 0.165977 0.124314 0.028412 0.021885 0.234008
MLP 0.010331 0.011387 0.010331 0.010000 0.024026 0.013581 5.913928
KNN 0.015266 0.00962 0.015266 0.014351 0.016323 0.006342 2.590869
RF 0.007615 0.010725 0.007615 0.008411 0.016425 0.007389 0.433674
LR 0.006048 0.005895 0.006048 0.006385 0.01425 0.005388 2.752432
DT 0.004937 0.005132 0.004937 0.004463 0.003758 0.002974 0.437479
Spike2Vec SVM 0.01981 0.02825 0.01981 0.02270 0.01555 0.00959 0.47610
NB 0.00864 0.03391 0.00864 0.01524 0.01116 0.00755 0.00741
MLP 0.00267 0.00347 0.00267 0.00356 0.00022 0.00257 5.75725
KNN 0.01714 0.02418 0.01714 0.01927 0.01990 0.01220 0.00993
RF 0.00814 0.00652 0.00814 0.00645 0.00206 0.00159 0.03782
LR 0.00726 0.01165 0.00726 0.01068 0.00633 0.00329 0.03185
DT 0.01457 0.01302 0.01457 0.01389 0.01716 0.00889 0.00797
PWM2Vec SVM 0.01386 0.01648 0.01386 0.01651 0.01519 0.00998 0.39967
NB 0.01718 0.02313 0.01718 0.02055 0.01788 0.00860 0.00820
MLP 0.01803 0.01902 0.01803 0.01891 0.01228 0.00753 7.02375
KNN 0.01043 0.01275 0.01043 0.01042 0.00859 0.00419 0.03936
RF 0.02213 0.01162 0.02213 0.02231 0.01838 0.01264 0.04062
LR 0.01802 0.01861 0.01802 0.02047 0.01716 0.01037 0.00798
DT 0.00925 0.01388 0.00925 0.01112 0.00742 0.00530 0.01723
String Kernel SVM 0.00952 0.00581 0.00952 0.00786 0.00160 0.01255 0.08849
NB 0.03677 0.05084 0.03677 0.03489 0.02923 0.01730 0.00058
MLP 0.02344 0.06729 0.02344 0.03073 0.03126 0.01631 4.14838
KNN 0.01650 0.01933 0.01650 0.01870 0.01455 0.00640 0.00274
RF 0.02286 0.03968 0.02286 0.02447 0.02789 0.01317 0.07689
LR 0.00959 0.03215 0.00959 0.01878 0.02042 0.00588 0.00183
DT 0.03106 0.03357 0.03106 0.03238 0.03658 0.02079 0.00885
WDGRL SVM 0.005067 0.00313 0.005067 0.004219 0.000845 0.007464 0.036333
NB 0.018709 0.028927 0.018709 0.012416 0.009551 0.007475 0.000338
MLP 0.00703 0.037114 0.00703 0.01059 0.011773 0.006923 0.932333
KNN 0.00703 0.005878 0.00703 0.006132 0.003794 0.002335 0.001434
RF 0.008417 0.011911 0.008417 0.008514 0.006176 0.003087 0.044674
LR 0.005067 0.016623 0.005067 0.011562 0.013997 0.004144 0.000185
DT 0.017918 0.01796 0.017918 0.018219 0.016001 0.009088 0.002927
Auto-Encoder SVM 0.00920 0.01191 0.00920 0.01139 0.00167 0.00084 0.41194
NB 0.12364 0.07618 0.12364 0.09485 0.03558 0.01111 0.04840
MLP 0.01107 0.01326 0.01107 0.01200 0.01462 0.00857 0.74852
KNN 0.01132 0.01190 0.01132 0.01157 0.01692 0.01040 0.02262
RF 0.00792 0.01075 0.00792 0.00992 0.01423 0.00511 0.36076
LR 0.00877 0.01212 0.00877 0.01124 0.00138 0.00081 0.30021
DT 0.00890 0.01560 0.00890 0.01162 0.02010 0.00778 0.21743
SeqVec SVM 0.00921 0.00812 0.00921 0.00927 0.00116 0.00061 0.25855
NB 0.10878 0.04510 0.10878 0.07567 0.03016 0.01570 0.04303
MLP 0.00943 0.01155 0.00943 0.01044 0.02354 0.01569 0.77613
KNN 0.01221 0.01064 0.01221 0.00991 0.01766 0.00902 0.02768
RF 0.00694 0.01022 0.00694 0.00697 0.01370 0.00901 0.33069
LR 0.01052 0.00901 0.01052 0.01049 0.00148 0.00071 0.10309
DT 0.00974 0.00994 0.00974 0.00919 0.02002 0.01160 0.07449
Protein Bert _ 0.008258 0.005042 0.008258 0.006823 0.001385 0.010888 0.076791
CCP SVM 0.00648 0.00599 0.00498 0.00745 0.01347 0.00793 0.32149
NB 0.01322 0.01478 0.01395 0.01544 0.01815 0.01133 0.53631
MLP 0.00698 0.00894 0.00643 0.00845 0.01322 0.00813 3.13462
KNN 0.01231 0.01056 0.01487 0.01254 0.01998 0.01676 0.76152
RF 0.00542 0.00854 0.00322 0.00233 0.01543 0.00953 0.91258
LR 0.00435 0.00743 0.00434 0.00643 0.00734 0.00532 0.26145
DT 0.00543 0.00512 0.00743 0.00832 0.01843 0.00743 0.29754
CCP-NN SVM 0.00499 0.00674 0.00612 0.00687 0.01611 0.00974 0.22764
NB 0.01527 0.01412 0.01357 0.01314 0.01809 0.01167 0.51931
MLP 0.00925 0.00814 0.00853 0.00847 0.01363 0.00874 1.49651
KNN 0.01156 0.01013 0.01432 0.01216 0.03311 0.01356 0.54301
RF 0.00567 0.00854 0.00565 0.00542 0.01653 0.00756 0.78123
LR 0.00654 0.00432 0.00632 0.00425 0.00753 0.00594 0.34325
DT 0.00712 0.00578 0.00713 0.00835 0.01831 0.00845 0.29543
TABLE VII: Classification results (standard deviation values over $5$ runs) on Protein Subcellular datasets for different evaluation metrics.
Detailed results along with their discussion of the Coronavirus Host and Human DNA datasets are reported in Section V-B and Section V-C, respectively. The observed improvement in classification results when using the proposed method over the original CCP method can be attributed to several technical and logical factors discussed in detail in Section V-F.
V-B Results For Coronavirus Host Data
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Coronavirus Host dataset are shown in Table VIII. The best values for each embedding method are underlined while overall best values among all methods are shown in bold. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics (other than classifier training runtime) and achieves a near-perfect predictive classification performance. Although the performance gain for CCP-NN compared to the best CCP-based results is not significant, however, it still outperforms CCP for all evaluation metrics other than training runtime.
The comparison of the best performing proposed method from Table VIII i.e. CCP-NN with Spike2Vec, with the existing baseline models is shown in Table IX for the Coronavirus Host dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Spike2Vec embedding achieves $1.7\%$ improvement compared to the second-best results (i.e. original Spike2Vec with Random Forest and Logistic Regression classifiers) and achieves a higher average accuracy score in the case of the Coronavirus Host dataset.
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
CCP ($\phi_{CCP}$) OHE SVM 0.819 0.817 0.819 0.815 0.651 0.820 45.784
NB 0.616 0.738 0.616 0.620 0.442 0.753 0.312
MLP 0.830 0.823 0.830 0.815 0.587 0.798 29.640
KNN 0.814 0.815 0.814 0.805 0.621 0.805 0.503
RF 0.846 0.843 0.846 0.835 0.667 0.828 1.185
LR 0.833 0.813 0.833 0.807 0.600 0.804 20.693
DT 0.845 0.847 0.845 0.837 0.677 0.835 0.283
Spike2Vec SVM 0.794 0.811 0.794 0.781 0.666 0.827 22.888
NB 0.606 0.742 0.606 0.557 0.442 0.729 1.609
MLP 0.829 0.840 0.829 0.822 0.649 0.835 416.592
KNN 0.805 0.817 0.805 0.802 0.626 0.818 2.226
RF 0.852 0.853 0.852 0.846 0.712 0.840 8.844
LR 0.769 0.796 0.769 0.757 0.630 0.796 60.627
DT 0.826 0.827 0.826 0.821 0.602 0.813 2.207
PWM2Vec SVM 0.820 0.816 0.820 0.811 0.643 0.846 4.292
NB 0.434 0.434 0.434 0.358 0.358 0.714 0.412
MLP 0.804 0.804 0.804 0.795 0.589 0.797 7.458
KNN 0.798 0.800 0.798 0.792 0.613 0.811 0.234
RF 0.833 0.829 0.833 0.825 0.671 0.845 6.868
LR 0.810 0.808 0.810 0.799 0.579 0.787 16.907
DT 0.793 0.790 0.793 0.789 0.616 0.808 2.642
Autoencoder SVM 0.630 0.620 0.630 0.605 0.234 0.615 3.189
NB 0.440 0.489 0.440 0.382 0.354 0.695 0.138
MLP 0.706 0.680 0.706 0.689 0.325 0.655 9.281
KNN 0.730 0.749 0.730 0.732 0.482 0.753 0.186
RF 0.818 0.819 0.818 0.810 0.646 0.799 7.642
LR 0.625 0.625 0.625 0.598 0.229 0.605 3.425
DT 0.755 0.753 0.755 0.750 0.516 0.764 1.218
CCP NN ($\phi_{CCP\_NN}$) OHE SVM 0.815 0.812 0.815 0.814 0.640 0.812 39.413
NB 0.596 0.711 0.596 0.597 0.354 0.704 0.295
MLP 0.830 0.806 0.830 0.808 0.575 0.796 31.610
KNN 0.813 0.814 0.813 0.802 0.592 0.785 0.508
RF 0.843 0.832 0.843 0.825 0.641 0.819 1.134
LR 0.778 0.763 0.778 0.750 0.393 0.700 20.612
DT 0.835 0.828 0.835 0.820 0.633 0.818 0.179
Spike2Vec SVM 0.827 0.832 0.827 0.818 0.758 0.871 17.080
NB 0.583 0.724 0.583 0.524 0.454 0.729 1.523
MLP 0.837 0.844 0.837 0.831 0.710 0.866 334.025
KNN 0.810 0.812 0.810 0.806 0.677 0.848 1.489
RF 0.855 0.855 0.855 0.849 0.770 0.872 5.547
LR 0.800 0.813 0.800 0.792 0.738 0.861 45.502
DT 0.845 0.846 0.845 0.841 0.703 0.864 1.343
PWM2Vec SVM 0.839 0.836 0.839 0.831 0.612 0.797 4.404
NB 0.648 0.700 0.648 0.633 0.476 0.746 0.390
MLP 0.825 0.826 0.825 0.818 0.584 0.788 6.728
KNN 0.781 0.787 0.781 0.777 0.591 0.781 0.237
RF 0.854 0.856 0.854 0.849 0.682 0.835 6.202
LR 0.824 0.815 0.824 0.811 0.545 0.755 15.188
DT 0.824 0.827 0.824 0.820 0.594 0.795 2.739
Autoencoder SVM 0.494 0.347 0.494 0.404 0.110 0.547 3.865
NB 0.468 0.578 0.468 0.490 0.153 0.647 0.141
MLP 0.696 0.671 0.696 0.671 0.283 0.631 9.500
KNN 0.729 0.733 0.729 0.721 0.447 0.715 0.168
RF 0.834 0.836 0.834 0.830 0.582 0.799 6.296
LR 0.464 0.332 0.464 0.372 0.099 0.538 2.518
DT 0.805 0.809 0.805 0.802 0.532 0.783 1.104
TABLE VIII: Classification results (averaged over $5$ runs) on Coronavirus Host dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
OHE SVM 0.822 0.834 0.822 0.823 0.728 0.839 389.128
NB 0.677 0.808 0.677 0.654 0.517 0.815 56.741
MLP 0.779 0.761 0.779 0.757 0.622 0.715 390.289
KNN 0.805 0.794 0.805 0.792 0.674 0.781 16.211
RF 0.836 0.831 0.836 0.822 0.709 0.832 151.911
LR 0.835 0.849 0.835 0.824 0.734 0.832 48.786
DT 0.824 0.833 0.824 0.811 0.679 0.810 21.581
Spike2Vec SVM 0.848 0.852 0.848 0.842 0.739 0.883 191.066
NB 0.661 0.768 0.661 0.661 0.522 0.764 10.220
MLP 0.815 0.837 0.815 0.814 0.640 0.835 46.624
KNN 0.782 0.794 0.782 0.781 0.686 0.832 82.112
RF 0.853 0.848 0.853 0.845 0.717 0.864 15.915
LR 0.853 0.852 0.853 0.846 0.757 0.879 60.620
DT 0.829 0.827 0.829 0.825 0.696 0.855 4.261
PWM2Vec SVM 0.799 0.806 0.799 0.801 0.648 0.859 44.793
NB 0.381 0.584 0.381 0.358 0.400 0.683 2.494
MLP 0.782 0.792 0.782 0.778 0.693 0.848 21.191
KNN 0.786 0.782 0.786 0.779 0.679 0.838 12.933
RF 0.836 0.839 0.836 0.828 0.739 0.862 7.690
LR 0.809 0.815 0.809 0.800 0.728 0.852 274.91
DT 0.801 0.802 0.801 0.797 0.633 0.829 4.537
Autoencoder SVM 0.602 0.588 0.602 0.590 0.519 0.759 2575.9
NB 0.261 0.520 0.261 0.303 0.294 0.673 21.74
MLP 0.486 0.459 0.486 0.458 0.216 0.594 29.93
KNN 0.763 0.764 0.763 0.755 0.547 0.784 18.51
RF 0.800 0.796 0.800 0.791 0.648 0.815 57.90
LR 0.717 0.750 0.717 0.702 0.564 0.812 11072.6
DT 0.772 0.767 0.772 0.765 0.571 0.808 121.36
String Kernel SVM 0.601 0.673 0.601 0.602 0.325 0.624 5.198
NB 0.230 0.665 0.230 0.295 0.162 0.625 0.131
MLP 0.647 0.696 0.647 0.641 0.302 0.628 42.322
KNN 0.613 0.623 0.613 0.612 0.310 0.629 0.434
RF 0.668 0.692 0.668 0.663 0.360 0.658 4.541
LR 0.554 0.724 0.554 0.505 0.193 0.568 5.096
DT 0.646 0.674 0.646 0.643 0.345 0.653 1.561
WDGRL SVM 0.329 0.108 0.329 0.163 0.029 0.500 2.859
NB 0.004 0.095 0.004 0.007 0.002 0.496 0.008
MLP 0.328 0.136 0.328 0.170 0.032 0.499 5.905
KNN 0.235 0.198 0.235 0.211 0.058 0.499 0.081
RF 0.261 0.196 0.261 0.216 0.051 0.499 1.288
LR 0.332 0.149 0.332 0.177 0.034 0.500 0.365
DT 0.237 0.202 0.237 0.211 0.054 0.498 0.026
SeqVec SVM 0.711 0.745 0.711 0.698 0.497 0.747 0.751
NB 0.503 0.636 0.503 0.554 0.413 0.648 0.012
MLP 0.718 0.748 0.718 0.708 0.407 0.706 10.191
KNN 0.815 0.806 0.815 0.809 0.588 0.800 0.418
RF 0.833 0.824 0.833 0.828 0.678 0.839 1.753
LR 0.673 0.683 0.673 0.654 0.332 0.660 1.177
DT 0.778 0.786 0.778 0.781 0.618 0.825 0.160
Protein Bert _ 0.799 0.806 0.799 0.789 0.715 0.841 15742.9
TAPE SVM 0.818 0.823 0.818 0.811 0.711 0.854 3.201
NB 0.482 0.587 0.482 0.442 0.400 0.712 0.494
MLP 0.812 0.819 0.812 0.802 0.665 0.828 3.737
KNN 0.793 0.797 0.793 0.789 0.633 0.818 0.150
RF 0.830 0.834 0.830 0.823 0.725 0.846 13.656
LR 0.779 0.797 0.779 0.764 0.628 0.794 11.325
DT 0.785 0.786 0.785 0.782 0.578 0.798 4.675
$\phi_{CCP\_NN}$ (ours) - Spike2Vec SVM 0.309 0.216 0.309 0.152 0.078 0.503 7.596
NB 0.172 0.392 0.172 0.107 0.116 0.523 0.174
MLP 0.341 0.324 0.341 0.272 0.201 0.543 218.782
KNN 0.419 0.412 0.419 0.412 0.361 0.629 1.140
RF 0.870 0.870 0.870 0.870 0.864 0.921 19.982
LR 0.309 0.280 0.309 0.152 0.077 0.503 1.911
DT 0.808 0.808 0.808 0.807 0.796 0.880 3.766
TABLE IX: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e., CCP-NN with Spike2Vec) with baselines on Coronavirus Host dataset for different evaluation metrics. The best values are in bold.
Table X displays the standard deviation (SD) outcomes (averaged over 5 runs) for both the baseline methods and proposed approach on the Coronavirus Host dataset. The results indicate that, in most cases, the SD values are relatively low, usually below 0.02. This observation suggests that the classification results remain consistent across different experimental runs with random train-test splits. The low variability in the SD values reflects the stability of the reported classification performance for both the proposed and baseline models.
Embeddings Algo. Acc. Prec. Recall F1 (Weig.) F1 (Macro) ROC AUC Train Time (sec.)
OHE SVM 0.010697 0.010309 0.010697 0.009847 0.006786 0.004762 1.818067
NB 0.014762 0.011664 0.014762 0.012998 0.008811 0.00592 0.047871
MLP 0.01903 0.027088 0.01903 0.022536 0.017811 0.006836 1.241431
KNN 0.005715 0.007481 0.005715 0.004432 0.005243 0.002815 0.405597
RF 0.011174 0.010344 0.011174 0.011743 0.013327 0.006616 0.201381
LR 0.059575 0.039483 0.059575 0.057232 0.060847 0.036954 96.708299
DT 0.010625 0.010962 0.010625 0.010682 0.011695 0.005174 0.157993
Spike2Vec SVM 0.02187 0.03118 0.02187 0.02506 0.01717 0.01059 0.52562
NB 0.00954 0.03743 0.00954 0.01682 0.01232 0.00833 0.00818
MLP 0.00295 0.00383 0.00295 0.00393 0.00025 0.00284 6.35600
KNN 0.01892 0.02670 0.01892 0.02128 0.02197 0.01347 0.01097
RF 0.00898 0.00720 0.00898 0.00712 0.00227 0.00176 0.04175
LR 0.00802 0.01287 0.00802 0.01179 0.00698 0.00363 0.03516
DT 0.01608 0.01437 0.01608 0.01533 0.01894 0.00981 0.00880
PWM2Vec SVM 0.01459 0.01735 0.01459 0.01737 0.01599 0.01051 0.42070
NB 0.01808 0.02434 0.01808 0.02163 0.01882 0.00905 0.00863
MLP 0.01898 0.02002 0.01898 0.01990 0.01293 0.00793 7.39342
KNN 0.01098 0.01342 0.01098 0.01097 0.00904 0.00441 0.04144
RF 0.02330 0.01223 0.02330 0.02348 0.01934 0.01331 0.04276
LR 0.01896 0.01959 0.01896 0.02155 0.01807 0.01092 0.00840
DT 0.00974 0.01461 0.00974 0.01171 0.00781 0.00558 0.01814
String Kernel SVM 0.00892 0.00545 0.00892 0.00737 0.00150 0.01176 0.08293
NB 0.03446 0.04765 0.03446 0.03270 0.02739 0.01621 0.00054
MLP 0.02197 0.06306 0.02197 0.02880 0.02930 0.01528 3.28975
KNN 0.01546 0.01811 0.01546 0.01752 0.01364 0.00600 0.00257
RF 0.02143 0.03719 0.02143 0.02293 0.02613 0.01234 0.07206
LR 0.00898 0.03013 0.00898 0.01760 0.01914 0.00551 0.00171
DT 0.02911 0.03146 0.02911 0.03035 0.03428 0.01948 0.00829
WDGRL SVM 0.008378 0.005078 0.008378 0.006888 0.00141 0.002417 0.034498
NB 0.008720 0.055455 0.00872 0.022475 0.022506 0.007747 0.000352
MLP 0.016103 0.010655 0.016103 0.018511 0.014246 0.007622 2.437331
KNN 0.010047 0.009635 0.010047 0.010275 0.011106 0.006126 0.002751
RF 0.013395 0.019497 0.013395 0.015266 0.018384 0.008316 0.035652
LR 0.007675 0.094971 0.007675 0.008903 0.005274 0.00175 0.001188
DT 0.009280 0.008941 0.00928 0.009266 0.007579 0.004472 0.004392
Autoencoder SVM 0.00956 0.00974 0.00956 0.01059 0.00127 0.00094 0.35010
NB 0.05871 0.04630 0.05871 0.05414 0.02578 0.01233 0.03049
MLP 0.00846 0.01142 0.00846 0.00776 0.01468 0.00882 1.00276
KNN 0.00631 0.00803 0.00631 0.00807 0.00699 0.00493 0.01380
RF 0.00338 0.00548 0.00338 0.00381 0.01029 0.00786 0.72100
LR 0.00982 0.00982 0.00982 0.01072 0.00128 0.00093 0.19975
DT 0.01025 0.00968 0.01025 0.00968 0.02163 0.00821 0.09998
SeqVec SVM 0.00729 0.00924 0.00729 0.00924 0.00102 0.00031 0.29267
NB 0.14408 0.06203 0.14408 0.11723 0.02721 0.01650 0.01298
MLP 0.01185 0.01247 0.01185 0.01129 0.02103 0.00999 0.68591
KNN 0.01281 0.01485 0.01281 0.01456 0.02329 0.01131 0.05557
RF 0.01050 0.01599 0.01050 0.01294 0.01490 0.00771 0.36725
LR 0.00795 0.00984 0.00795 0.01007 0.00126 0.00053 0.22741
DT 0.01119 0.01183 0.01119 0.01277 0.02537 0.00989 0.12363
Protein Bert _ 0.02965 0.03204 0.02965 0.03091 0.03492 0.01984 0.00845
$\phi_{CCP\_NN}$ (ours) - Spike2Vec SVM 0.011316 0.020199 0.011316 0.017867 0.026025 0.014217 0.279999
NB 0.012632 0.018465 0.012632 0.012929 0.013448 0.006475 0.009835
MLP 0.007226 0.009827 0.007226 0.007801 0.014985 0.009328 4.298934
KNN 0.010121 0.012496 0.010121 0.0117 0.024522 0.012687 1.301837
RF 0.007323 0.011286 0.007323 0.00848 0.006726 0.001826 0.159344
LR 0.014064 0.02447 0.014064 0.016745 0.010772 0.005139 0.430449
DT 0.006303 0.011474 0.006303 0.007942 0.012572 0.006275 0.016553
TABLE X: Standard Deviation values of 5 runs for Classification results on the proposed and baseline methods for Coronavirus Host dataset.
V-C Results For Human DNA Data
Table XI presents the classification results for both the proposed CCP-NN and the conventional CCP approach on the Human DNA dataset. The results are averaged over 5 runs. The best values for each embedding method are underlined, and the overall best values among all methods are shown in bold. It is evident that the proposed CCP-NN consistently outperforms the original CCP-based low-dimensional representation for all evaluation metrics, except for the classifier training runtime. The classification accuracy achieved by CCP-NN is notably higher, with the best-performing results using CCP-NN (with Autoencoder and Random Forest Classifier) showing a significant improvement of $10.8\%$ compared to the best results obtained from the original CCP (using One-Hot Encoding with Random Forest classifier).
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
CCP ($\phi_{CCP}$) OHE SVM 0.370 0.466 0.370 0.260 0.213 0.553 6.585
NB 0.375 0.598 0.375 0.371 0.369 0.634 0.138
MLP 0.667 0.672 0.667 0.669 0.644 0.795 5.791
KNN 0.581 0.595 0.581 0.580 0.549 0.732 0.189
RF 0.762 0.813 0.762 0.763 0.763 0.829 1.789
LR 0.378 0.541 0.378 0.270 0.228 0.559 4.224
DT 0.661 0.663 0.661 0.661 0.638 0.791 0.554
Spike2Vec SVM 0.447 0.417 0.447 0.385 0.303 0.602 1.212
NB 0.215 0.313 0.215 0.180 0.148 0.543 0.012
MLP 0.526 0.523 0.526 0.516 0.467 0.687 10.664
KNN 0.601 0.612 0.601 0.602 0.565 0.750 0.137
RF 0.725 0.753 0.725 0.721 0.709 0.806 2.693
LR 0.437 0.443 0.437 0.385 0.312 0.605 0.184
DT 0.593 0.597 0.593 0.594 0.561 0.746 0.142
PWM2Vec SVM 0.312 0.302 0.312 0.162 0.085 0.505 2.884
NB 0.095 0.324 0.095 0.058 0.051 0.508 0.017
MLP 0.311 0.312 0.311 0.165 0.088 0.505 318.449
KNN 0.194 0.320 0.194 0.113 0.080 0.509 1.537
RF 0.315 0.346 0.315 0.178 0.104 0.509 1.952
LR 0.313 0.277 0.313 0.166 0.089 0.506 0.826
DT 0.310 0.293 0.310 0.174 0.100 0.507 0.062
Autoencoder SVM 0.440 0.504 0.440 0.394 0.341 0.606 3.620
NB 0.192 0.312 0.192 0.147 0.147 0.539 0.076
MLP 0.485 0.483 0.485 0.481 0.440 0.672 33.924
KNN 0.496 0.498 0.496 0.494 0.462 0.685 0.262
RF 0.593 0.715 0.593 0.585 0.576 0.713 14.667
LR 0.414 0.452 0.414 0.370 0.308 0.593 2.159
DT 0.476 0.479 0.476 0.476 0.447 0.679 3.027
CCP Nearest Neighbor ($\phi_{CCP\_NN}$) OHE SVM 0.618 0.621 0.618 0.614 0.588 0.749 5.475
NB 0.357 0.573 0.357 0.344 0.343 0.624 0.141
MLP 0.667 0.669 0.667 0.666 0.639 0.791 5.556
KNN 0.569 0.581 0.569 0.568 0.538 0.729 0.178
RF 0.768 0.828 0.768 0.771 0.776 0.834 1.727
LR 0.579 0.595 0.579 0.569 0.536 0.713 5.182
DT 0.675 0.677 0.675 0.675 0.653 0.801 0.498
Spike2Vec SVM 0.309 0.226 0.309 0.153 0.087 0.506 1.467
NB 0.204 0.280 0.204 0.163 0.169 0.533 0.013
MLP 0.307 0.296 0.307 0.283 0.236 0.554 30.027
KNN 0.361 0.360 0.361 0.357 0.314 0.604 0.130
RF 0.558 0.690 0.558 0.545 0.530 0.688 3.561
LR 0.322 0.388 0.322 0.193 0.133 0.519 0.216
DT 0.532 0.532 0.532 0.531 0.502 0.709 0.234
PWM2Vec SVM 0.316 0.161 0.316 0.160 0.080 0.503 2.380
NB 0.069 0.301 0.069 0.032 0.038 0.506 0.015
MLP 0.314 0.147 0.314 0.158 0.078 0.502 131.941
KNN 0.222 0.303 0.222 0.106 0.067 0.504 1.198
RF 0.315 0.190 0.315 0.161 0.081 0.503 0.941
LR 0.314 0.140 0.314 0.159 0.079 0.502 0.494
DT 0.315 0.192 0.315 0.163 0.083 0.503 0.005
Autoencoder SVM 0.309 0.216 0.309 0.152 0.078 0.503 7.596
NB 0.172 0.392 0.172 0.107 0.116 0.523 0.174
MLP 0.341 0.324 0.341 0.272 0.201 0.543 218.782
KNN 0.419 0.412 0.419 0.412 0.361 0.629 1.140
RF 0.870 0.870 0.870 0.870 0.864 0.921 19.982
LR 0.309 0.280 0.309 0.152 0.077 0.503 1.911
DT 0.808 0.808 0.808 0.807 0.796 0.880 3.766
TABLE XI: Classification results (averaged over $5$ runs) on Human DNA dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
Embeddings Algo. Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F1 (Weig.) $\uparrow$ F1 (Macro) $\uparrow$ ROC AUC $\uparrow$ Train Time (sec.) $\downarrow$
OHE SVM 0.579 0.599 0.579 0.576 0.561 0.721 10.475
NB 0.165 0.142 0.165 0.101 0.125 0.529 0.746
MLP 0.600 0.611 0.600 0.612 0.564 0.723 45.785
KNN 0.638 0.649 0.638 0.640 0.598 0.754 1.574
RF 0.722 0.768 0.722 0.741 0.729 0.811 5.749
LR 0.566 0.568 0.566 0.574 0.521 0.698 9.781
DT 0.611 0.615 0.611 0.619 0.590 0.747 0.749
Spike2Vec SVM 0.597 0.602 0.597 0.589 0.563 0.733 4.612
NB 0.175 0.143 0.175 0.106 0.128 0.532 0.039
MLP 0.618 0.618 0.618 0.613 0.573 0.747 22.292
KNN 0.640 0.653 0.640 0.642 0.608 0.772 0.561
RF 0.752 0.773 0.752 0.749 0.736 0.824 2.558
LR 0.569 0.570 0.569 0.555 0.525 0.710 2.074
DT 0.621 0.624 0.621 0.621 0.594 0.765 0.275
PWM2Vec SVM 0.302 0.241 0.302 0.165 0.091 0.505 10011.3
NB 0.084 0.442 0.084 0.063 0.066 0.511 4.565
MLP 0.310 0.350 0.310 0.175 0.107 0.510 320.555
KNN 0.121 0.337 0.121 0.093 0.077 0.509 2.193
RF 0.309 0.332 0.309 0.181 0.110 0.510 65.250
LR 0.304 0.257 0.304 0.167 0.094 0.506 23.651
DT 0.306 0.284 0.306 0.181 0.111 0.509 1.861
Autoencoder SVM 0.621 0.638 0.621 0.624 0.593 0.769 22.230
NB 0.260 0.426 0.260 0.247 0.268 0.583 0.287
MLP 0.621 0.624 0.621 0.620 0.578 0.756 111.809
KNN 0.565 0.577 0.565 0.568 0.547 0.732 1.208
RF 0.689 0.738 0.689 0.683 0.668 0.774 20.131
LR 0.692 0.700 0.692 0.693 0.672 0.799 58.369
DT 0.543 0.546 0.543 0.543 0.515 0.718 10.616
String Kernel SVM 0.618 0.617 0.618 0.613 0.588 0.753 39.791
NB 0.338 0.452 0.338 0.347 0.333 0.617 0.276
MLP 0.597 0.595 0.597 0.593 0.549 0.737 331.068
KNN 0.645 0.657 0.645 0.646 0.612 0.774 1.274
RF 0.731 0.776 0.731 0.729 0.723 0.808 12.673
LR 0.571 0.570 0.571 0.558 0.532 0.716 2.995
DT 0.630 0.631 0.630 0.630 0.598 0.767 2.682
WDGRL SVM 0.318 0.101 0.318 0.154 0.069 0.500 0.751
NB 0.232 0.214 0.232 0.196 0.138 0.517 0.004
MLP 0.326 0.286 0.326 0.263 0.186 0.535 8.613
KNN 0.317 0.317 0.317 0.315 0.266 0.574 0.092
RF 0.453 0.501 0.453 0.430 0.389 0.625 1.124
LR 0.323 0.279 0.323 0.177 0.095 0.507 0.041
DT 0.368 0.372 0.368 0.369 0.328 0.610 0.047
SeqVec SVM 0.656 0.661 0.656 0.652 0.611 0.791 0.891
NB 0.324 0.445 0.312 0.295 0.282 0.624 0.036
MLP 0.657 0.633 0.653 0.646 0.616 0.783 12.432
KNN 0.592 0.606 0.592 0.591 0.552 0.717 0.571
RF 0.713 0.724 0.701 0.702 0.693 0.752 2.164
LR 0.725 0.715 0.726 0.725 0.685 0.784 1.209
DT 0.586 0.553 0.585 0.577 0.557 0.736 0.24
Protein Bert _ 0.542 0.580 0.542 0.514 0.447 0.675 58681.57
TAPE SVM 0.601 0.600 0.601 0.598 0.567 0.748 18.323
NB 0.231 0.334 0.231 0.224 0.213 0.558 0.340
MLP 0.538 0.533 0.538 0.526 0.462 0.689 7.882
KNN 0.532 0.539 0.532 0.533 0.503 0.712 0.173
RF 0.689 0.705 0.689 0.686 0.672 0.791 42.389
LR 0.544 0.548 0.544 0.519 0.456 0.683 19.215
DT 0.582 0.585 0.582 0.582 0.553 0.743 15.769
$\phi_{CCP\_NN}$ (ours) - Autoencoder SVM 0.309 0.216 0.309 0.152 0.078 0.503 7.596
NB 0.172 0.392 0.172 0.107 0.116 0.523 0.174
MLP 0.341 0.324 0.341 0.272 0.201 0.543 218.782
KNN 0.419 0.412 0.419 0.412 0.361 0.629 1.140
RF 0.870 0.870 0.870 0.870 0.864 0.921 19.982
LR 0.309 0.280 0.309 0.152 0.077 0.503 1.911
DT 0.808 0.808 0.808 0.807 0.796 0.880 3.766
TABLE XII: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e. CCP-NN with Autoencoder) with baselines on Human DNA dataset for different evaluation metrics. The best values are in bold.
The comparison of the best performing proposed method from Table XI, i.e. CCP-NN with Autoencoder, with the existing baseline models is shown in Table XII for the Human DNA dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Autoencoder embedding achieves $11.8\%$ improvement compared to the second-best results (i.e. original Spike2Vec with Random Forest classifier).
It is noteworthy that the pre-trained Protein Bert exhibited significantly poorer performance on the Human DNA dataset compared to its performance on the Protein Subcellular and Coronavirus Host datasets. The underlying reason for this discrepancy lies in the fact that the Protein Bert model is designed and trained specifically on molecular sequence data. Consequently, when faced with nucleotide sequences of Human DNA, the model struggles to generalize effectively, leading to its subpar performance. In contrast, the proposed method demonstrated the highest performance among all approaches, outperforming the baseline methods on the Human DNA dataset. This indicates the robustness and efficacy of our proposed method in handling diverse biological sequence data.
Table XIII presents the standard deviation (SD) results (averaged over 5 runs) for the baseline methods and our proposed approach to the Human DNA dataset. The findings reveal that the majority of the standard deviations. values are relatively low, generally below 0.02. This observation indicates that the classification results exhibit consistency across various experimental runs with random train-test splits. The low variability in the SD values highlights the stability of the reported classification performance for both the proposed and baseline models.
Embeddings Algo. Acc. Prec. Recall F1 (Weig.) F1 (Macro) ROC AUC Train Time (sec.)
OHE SVM 0.009508 0.011373 0.009508 0.012585 0.011098 0.004223 0.669254
NB 0.137802 0.030246 0.137802 0.121778 0.019086 0.010169 0.023381
MLP 0.013534 0.017288 0.013534 0.01404 0.014754 0.00558 1.455186
KNN 0.00772 0.008617 0.00772 0.008645 0.024282 0.010772 0.215671
RF 0.005861 0.014158 0.005861 0.007086 0.012801 0.005187 0.316404
LR 0.009137 0.005479 0.009137 0.011615 0.003251 0.001162 0.125465
DT 0.008326 0.008945 0.008326 0.008091 0.007818 0.004461 0.100935
Spike2Vec SVM 0.02377 0.03390 0.02377 0.02724 0.01866 0.01151 0.57132
NB 0.01037 0.04069 0.01037 0.01829 0.01339 0.00906 0.00889
MLP 0.00321 0.00416 0.00321 0.00427 0.00027 0.00308 6.90869
KNN 0.02056 0.02902 0.02056 0.02313 0.02388 0.01464 0.01192
RF 0.00977 0.00783 0.00977 0.00774 0.00247 0.00191 0.04538
LR 0.00871 0.01398 0.00871 0.01282 0.00759 0.00395 0.03822
DT 0.01748 0.01562 0.01748 0.01666 0.02059 0.01067 0.00957
PWM2Vec SVM 0.01751 0.02082 0.01751 0.02085 0.01919 0.01261 0.50484
NB 0.02170 0.02921 0.02170 0.02596 0.02258 0.01086 0.01035
MLP 0.02278 0.02402 0.02278 0.02388 0.01552 0.00951 8.87211
KNN 0.01318 0.01611 0.01318 0.01316 0.01085 0.00529 0.04972
RF 0.02796 0.01468 0.02796 0.02818 0.02321 0.01597 0.05131
LR 0.02276 0.02351 0.02276 0.02586 0.02168 0.01310 0.01008
DT 0.01169 0.01753 0.01169 0.01405 0.00937 0.00669 0.02177
Autoencoder SVM 0.007347 0.031924 0.007347 0.007065 0.028499 0.013436 0.434526
NB 0.01508 0.019538 0.01508 0.015753 0.024613 0.009869 0.017177
MLP 0.009726 0.013115 0.009726 0.010498 0.021136 0.009565 1.394414
KNN 0.005735 0.005619 0.005735 0.006138 0.037517 0.016833 0.094046
RF 0.003454 0.004077 0.003454 0.003639 0.028805 0.01733 0.185307
LR 0.011574 0.032636 0.011574 0.013635 0.02211 0.008987 0.27776
DT 0.00641 0.008259 0.00641 0.006691 0.036235 0.008376 0.108608
String Kernel SVM 0.00981 0.00599 0.00981 0.00811 0.00165 0.01293 0.09123
NB 0.03791 0.05241 0.03791 0.03597 0.03013 0.01783 0.00059
MLP 0.02417 0.06937 0.02417 0.03168 0.03223 0.01681 4.27668
KNN 0.01701 0.01992 0.01701 0.01928 0.01500 0.00659 0.00283
RF 0.02357 0.04091 0.02357 0.02523 0.02875 0.01358 0.07927
LR 0.00988 0.03315 0.00988 0.01936 0.02105 0.00606 0.00188
DT 0.03202 0.03461 0.03202 0.03338 0.03771 0.02143 0.00912
WDGRL SVM 0.003191 0.001912 0.003191 0.002604 0.00054 0.003424 0.040458
NB 0.013202 0.015189 0.013202 0.017859 0.015814 0.007535 0.000162
MLP 0.013313 0.021275 0.013313 0.01608 0.015356 0.007229 2.113733
KNN 0.007286 0.010892 0.007286 0.010094 0.008831 0.003216 0.000944
RF 0.011424 0.022522 0.011424 0.012721 0.018023 0.008342 0.022051
LR 0.003251 0.011279 0.003251 0.004737 0.003721 0.000956 0.0014
DT 0.009034 0.011169 0.009034 0.00988 0.015741 0.00895 0.004752
SeqVec SVM 0.004323 0.0053 0.004323 0.002444 0.00354 0.002263 0.249421
NB 0.003381 0.007385 0.003381 0.003252 0.002496 0.002215 0.004853
MLP 0.006487 0.00829 0.006487 0.005008 0.003721 0.001173 1.397002
KNN 0.005323 0.005685 0.005323 0.004541 0.002293 0.001867 0.082383
RF 0.008888 0.009345 0.008888 0.007638 0.005948 0.003357 0.592595
LR 0.00786 0.003496 0.00786 0.006579 0.002844 0.002397 3.962773
DT 0.004853 0.006327 0.004853 0.005272 0.005007 0.002373 0.324555
Protein Bert _ 0.02685 0.02874 0.02547 0.02874 0.03024 0.01774 0.00788
TAPE SVM 0.00877 0.006212 0.00877 0.00792 0.005215 0.002805 0.514809
NB 0.020059 0.026968 0.020059 0.022348 0.018922 0.011747 0.0111
MLP 0.017366 0.020384 0.017366 0.018008 0.022814 0.010171 2.903999
KNN 0.006358 0.008292 0.006358 0.007089 0.009437 0.004215 0.031404
RF 0.00615 0.006748 0.00615 0.005848 0.009419 0.00505 5.711691
LR 0.009299 0.018944 0.009299 0.009405 0.010728 0.006394 1.191378
DT 0.01841 0.018301 0.01841 0.018729 0.016768 0.009301 1.218549
$\phi_{CCP\_NN}$ (ours) - Autoencoder SVM 0.009508 0.011373 0.009508 0.012585 0.011098 0.004223 0.669254
NB 0.137802 0.030246 0.137802 0.121778 0.019086 0.010169 0.023381
MLP 0.013534 0.017288 0.013534 0.01404 0.014754 0.00558 1.455186
KNN 0.00772 0.008617 0.00772 0.008645 0.024282 0.010772 0.215671
RF 0.005861 0.014158 0.005861 0.007086 0.012801 0.005187 0.316404
LR 0.009137 0.005479 0.009137 0.011615 0.003251 0.001162 0.125465
DT 0.008326 0.008945 0.008326 0.008091 0.007818 0.004461 0.100935
TABLE XIII: Standard Deviation values of 5 runs for Classification results on the proposed and baseline methods for the Human DNA dataset.
V-D Runtime Evaluation
[…]
where $R_{\phi_{CCP}}$ represents the runtime of the original CCP method while $R_{\phi_{CCP\_NN}}$ corresponds to the runtime for our NN-based CCP computation.
The computational runtime for $R_{\phi_{CCP\_NN}}$ and $R_{\phi_{CCP}}$ along with the performance gain is reported in Table XVI,  XVI, and  XVI for the Protein Subcellular, Human DNA, and Coronavirus Host datasets, respectively. For protein subcellular data, we can observe that for all $4$ embedding methods as input to the CCP and CCP-NN, the performance gain (i.e. Percentage improvement) for our CCP-NN is $23.32\%$, $72.34\%$, $55.93\%$, and $92.88\%$, for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. For the Coronavirus Host dataset, we can again observe that in terms of computations runtime performance gain, the proposed CCP-NN significantly outperforms the original CCP by $39.429\%$, $55.950\%$, $97.865\%$, and $93.989\%$ for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. Similarly, for the Human DNA dataset, we can observe $94.385\%$ (OHE), $90.243\%$ (for Spike2Vec), $91.456\%$ (PWM2Vec), and $85.425\%$ (for Autoencoder) improvement in computational runtime for CCP-NN compared to the CCP.
To observe the increase in runtime with the increasing number of data points (i.e. embeddings), we select the overall best-performing embedding method with CCP and CCP-NN, i.e. Autoencoder, and compute runtime with an increasing number of embeddings. The runtime results are reported in Figure 4. We can observe that the vanilla CCP’s runtime increases very quickly as we increase the number of embeddings. On the other hand, the runtime increase for the CCP-NN is very slow, showing its better scalability property.
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
OHE 1358.408 1772.998 23.32%
Spike2Vec 392.002 1417.402 72.34%
PWM2Vec 664.225 1507.067 55.93%
Autoencoder 74.364 1058.623 92.88%
TABLE XIV: Runtime comparison for CCP and CCP-NN using different embedding for Protein Subcellular Data.
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
OHE 2875.668 4747.570 39.429%
Spike2Vec 650.596 1476.935 55.950%
PWM2Vec 131.138 6141.813 97.865%
Autoencoder 44.404 738.684 93.989%
TABLE XV: The runtime comparison for CCP and CCP-NN using different embedding for Coronavirus Host Data.
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
OHE 13.640 242.940 94.385 %
Spike2Vec 27.544 282.286 90.243%
PWM2Vec 7.093 83.026 91.456 %
Autoencoder 166.852 1144.768 85.425%
TABLE XVI: The runtime comparison for CCP and CCP-NN using different embedding for Human DNA Data.
(a) Protein Subcellular
(b) Coronavirus Host
(c) Human DNA
Figure 4: Runtime for embedding generation of Autoencoder with an increasing number of data points for different datasets. The figure is best seen in color.
V-E Statistical Significance
We used the Student’s t-test and calculated the $p$-values using the average and standard deviations (SD) of $5$ runs to determine whether the computed classification results are statistically significant. Given that SD values are extremely low for all datasets, i.e. $<0.002$ in most cases, we noticed that $p$-values were $<0.05$ in the majority of cases and for all embedding methods, hence validating the statistical significance of the findings.
V-F Discussion
The observed improvement in classification results when using the proposed method over the original CCP method can be attributed to several technical and logical factors.
Efficient Nearest Neighbor Search
The proposed method employs the nearest neighbor (NN) search technique using the AnnoyIndex. By leveraging NN, the algorithm reduces the computational complexity of pairwise distance calculations. In high-dimensional spaces like biological sequencing data where feature dimensions can be large, the NN-based approach significantly speeds up the computation, allowing for better handling of the complexity and potentially better capturing of correlations between features.
Handling High-Dimensional Data
In sequence classification, feature spaces can often be high-dimensional due to the representation of various attributes. The proposed method can better preserve the feature information in the low-dimensional space than the original CCP due to efficient nearest-neighbor computation, leading to more meaningful density estimations. The nearest neighbor search focuses on capturing local patterns in data, which are particularly relevant for correlations between features. Using the density estimation from an NN-based neighbor search, the method effectively identifies relationships between relevant features, thereby representing the underlying patterns better.
Robustness to Noisy Data
The data might contain noise and outliers, which can negatively impact the accuracy of correlation estimation. The NN-based method can be more robust to noisy data as it focuses on local patterns rather than global distances. It implicitly handles noise using neighbor relationships, leading to more reliable density estimations and better classification results.
Optimal Feature Clustering
The proposed method partitions the features into clusters based on their correlation patterns. By employing NN-based density estimation, the method identifies more optimal feature clusters, which in turn can enhance the separation of target classes. The ability to detect relevant feature subsets for classification can contribute to improved accuracy.
In general, the superior classification results obtained with the proposed method can be attributed to its efficient handling of high-dimensional data, the use of NN for nearest neighbor search, enhanced correlation estimation, robustness to noise, and optimal feature clustering. These advantages collectively enable the method to capture the underlying patterns and provide more discriminative representations for improved classification performance as demonstrated in a variety of datasets including protein subcellular localization data, Coronavirus Host data, and Human DNA in our experiments.
VI Conclusion
In this study, we addressed the challenges of analyzing molecular sequence data, which involves a large number of sequences and complex protein structures. We proposed an efficient and fast method called Nearest Neighbor Correlated Clustering and Projection (CCP-NN). The CCP-NN method is based on the original Correlated Clustering and Projection (CCP) technique, but it incorporates an NN search for computing the density map and correlations. Through a series of experimental evaluations, we compared the performance of CCP and CCP-NN in classifying molecular sequences. The results demonstrated that CCP-NN outperforms CCP in terms of classification accuracy while also reducing computational runtime. Future work includes further enhancements to the CCP-NN framework, such as incorporating additional information or integrating it with other machine-learning methods for comprehensive analysis of sequences.